The least of three consecutive odd integers is 2 more than half the largest. What are the integers?
Answer by checkley79(3341)

LET X, X+1 & X+2 BE THE THREE CONSECUTIVE INTEGERS.
X=(X+2)/2+2
X-2=(X+2)/2 CROSS MULTIPLY.
2(X-2)=X+2
2X-4=X+2
2X-X=2+4
X=6 ANS. FOR THE SMALLEST INTEGER.
PROOF:
X+2=6+2=8 ANS. FOR THE LARGEST INTEGER.
6=(6+2)/2+2
6=8/2+2
6=4+2
6=6